OpenStudy (anonymous):

how do you solve [6y-(7x-2)][6y+(7x-2)]

OpenStudy (anonymous):

why do i square it?

OpenStudy (anonymous):

You can treat it like a binomial difference of squares (6y - a) (6y +a) where a = 7x-2

OpenStudy (anonymous):

what about the other side?

OpenStudy (anonymous):

I did wonder about "solve"...

OpenStudy (anonymous):

no it says find the product...

OpenStudy (anonymous):

nvm thats simplify...

OpenStudy (anonymous):

but still, do i do the same thing on the other side?

OpenStudy (anonymous):

what other side?

OpenStudy (anonymous):

its [6y-(7x-2)][6y+(7x-2)]?

OpenStudy (anonymous):

OK, so there is just that, no sides....

OpenStudy (anonymous):

and you want to multiply them together...

OpenStudy (anonymous):

i have to find the product, yes.

OpenStudy (anonymous):

Yes, that's the same thing. The first post from Ryan is a good answer. Are u expected to expand it all the way out.....?

OpenStudy (anonymous):

no i dont think so

OpenStudy (anonymous):

Then I would go with Ryan's answer.

OpenStudy (anonymous):

so just 6y-a 6y+a?

OpenStudy (anonymous):

Hmm... he appears to have deleted his answer..

OpenStudy (anonymous):

so what do i do?

OpenStudy (anonymous):

It was (6y)^2 - (7x-2)^2 which u could write as 36y^2 - (7x-2)^2 if u wanted to.

OpenStudy (anonymous):

oh , alright

OpenStudy (anonymous):

r u sure u are not expected to expand it all out?

OpenStudy (dumbcow):

\[(a+b)(a-b) = a^{2} - b^{2}\] a=6y b = 7x-2

OpenStudy (anonymous):

\[(a+b)(a-b)=a^2-b^2\] put \[a=6y, b=7x-2\] get \[(6y)^2-(72-2)^2\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, u r done, then.

OpenStudy (anonymous):

\[(6y)^2-(7x-2)^2\]

OpenStudy (anonymous):

satallite where did 72 come from?

OpenStudy (anonymous):

Typo

OpenStudy (anonymous):

you can multiply out if you like. get \[36y^2-(49x^2-28x+4)\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Now we are repeating......